# A Lightweight Method for Defending Against UAF Vulnerabilities

### v1

@@
Abstract.
The widespread presence of Use-After-Free (UAF) vulnerabilities poses a serious threat to software security, with dangling pointers being considered the primary cause of these vulnerabilities. However, existing methods for defending against UAF vulnerabilities by eliminating dangling pointers need to interrupt the program’s execution when encountering pointer assignment operations to look up the objects pointed to by the pointers and store the memory addresses of the pointers in a specific data structure. This makes these methods not lightweight. To overcome this drawback, we propose a novel approach called LightDE. This method does not require storing the memory addresses of pointers or locating the objects pointed to by pointers during program execution. LightDE uses our proposed structure-sensitive pointer analysis method to determine the objects pointed to by pointers and stores the pointing relationships in the program’s data segment during program compilation. Since LightDE only needs to check whether the pointers identified by the pointer analysis point to the released objects when the objects are released, LightDE is very lightweight. Our experimental results show that LightDE can effectively defend against UAF vulnerabilities, and the additional performance overhead it introduces is very low.
@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
@@
Use-After-Free (UAF) vulnerabilities pose a serious threat to program security as they can lead to arbitrary code execution, privilege escalation, and data leakage (Gui et al., 2021; Song et al., 2019; Szekeres et al., 2013). Unfortunately, the number of UAF vulnerabilities has been increasing over the years, and their impact has become increasingly severe (Gui et al., 2021; Yan et al., 2018; Cuoq et al., 2012; Ye et al., 2014). Table 1 shows the changes in the number of UAF vulnerabilities in the NVD database from 2011 to 2023, as well as the number of high-threat UAF vulnerabilities. There are two types of UAF vulnerabilities: stack UAF and heap UAF (Younan, 2015; Van Der Kouwe et al., 2018; Szekeres et al., 2013). Stack UAF occurs when a dangling pointer is used to access stack memory that has already been freed. However, modern compilers generate warning messages for such vulnerabilities, making it easy for programmers to detect and fix them. Additionally, stack space allocation and deallocation are automatic, making stack UAF vulnerabilities difficult to exploit. Heap UAF refers to accessing freed heap memory through a dangling pointer. The difficulty in defending against heap UAF primarily stems from the flexible allocation and deallocation of heap memory, as well as the complexity of pointer propagation. All UAF vulnerabilities discussed in this paper are heap UAF.
Figure 1. Use-after-free Vulnerabilities in NVD (NIST, 2005).
@@
Researchers have proposed various methods to defend against UAF vulnerabilities, but each approach has its limitations and drawbacks. One method involves checking the validity of pointer access during pointer dereferencing, which requires verifying memory access legality before each pointer dereference (Nagarakatte et al., 2010, 2012; Necula et al., 2002; Condit et al., 2003; Zhang et al., 2019). However, this method can lead to incompatibility with existing libraries and introduce unacceptable runtime overhead. An alternative approach modifies memory allocation and deallocation strategies, making it harder for attackers to exploit UAF vulnerabilities (Boehm et al., 2002; Oiwa, 2009; Zhang and Zonouz, 2018; Novark and Berger, 2010; Berger and Zorn, 2006; Dang et al., 2017). However, this method only provides probabilistic defense, and attackers can bypass it through techniques like heap spraying and other attack methods. Recent research has proposed defending against UAF vulnerabilities by eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020). Eliminating dangling pointers is a very promising method for mitigating UAF vulnerabilities, as dangling pointers are the root cause of these vulnerabilities. However, existing methods of eliminating dangling pointers need to record pointer memory addresses and look up the objects pointed to by pointers during program execution, making these methods not lightweight.
@@
Since dangling pointers are the root cause of UAF vulnerabilities (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020), this paper focuses on defending against UAF vulnerabilities by eliminating dangling pointers. Current methods (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020) for defending against UAF vulnerabilities by eliminating dangling pointers, rely on inserting functions after pointer assignment instructions during compilation to track pointers and the objects they reference. This method requires the program to pause its normal execution flow when encountering a pointer assignment instruction and to execute the logic of the inserted function. The role of the inserted code is to look up the objects pointed to by the pointers and record the memory addresses of the pointers in the object’s pointer set. Although existing dangling pointer elimination schemes use different methods to locate the objects pointed to by pointers and various data structures to store the memory addresses of pointers to improve program execution efficiency, the basic principle remains the same. In this process, locating the objects pointed to by pointers and recording the memory addresses of pointers are time-consuming, making current dangling pointer elimination methods not lightweight.
@@
To address the issue of existing methods being unable to eliminate dangling pointers in a lightweight manner, we propose LightDE, a lightweight approach to eliminate dangling pointers and defend against UAF vulnerabilities. LightDE’s uniqueness lies in its combination of pointer analysis (Lhotak et al., 2013; Hardekopf and Lin, 2007; Pereira and Berlin, 2009; Kroening and Tautschnig, 2014) to determine the objects pointed to by pointers at compile time, and storing the pointing relationships in the program’s data segment during compilation. The significant advantage of our method over existing methods is that it does not require locating the objects pointed to by pointers or storing the memory addresses of pointers during program execution. Additionally, we use an asynchronous method to eliminate dangling pointers by running a separate thread on an idle CPU core. This thread uses the pointing relationships stored in the program’s data segment to check if the pointers identified by the pointer analysis point to the released objects and eliminate the corresponding dangling pointers.
@@
We implemented a prototype system of LightDE and demonstrated its effectiveness using real UAF vulnerabilities. Our experimental results on the SPEC CPU2006 (Henning, 2006) benchmark suite show that LightDE introduces very low additional runtime overhead, with a geometric mean of only 8.66%. The geometric mean of the additional memory overhead introduced by LightDE is 7.26%. Experimental results on the SPLASH-2X (Group et al., 2011) benchmark suite show that LightDE scales well on multithreaded applications.
@@
1. We propose a lightweight method for eliminating dangling pointers, LightDE. Our approach determines pointer relationships during program compilation and embeds this information into the program’s data segment at compile time. This allows our method to avoid interrupting the program’s execution flow to locate pointer targets or store the memory addresses of pointers, making LightDE very lightweight.
2. We propose a structure-sensitive pointer analysis method to obtain the structural information of objects and improve the accuracy of pointer analysis results.
3. We developed a prototype system of LightDE based on LLVM 13 and SVF-2.5 to validate our design. We conducted an experiment to evaluate the LightDE prototype. The experimental results show that LightDE can effectively defend against UAF vulnerabilities while introducing very low additional performance overhead.
2. Motivation Example
1typedef struct {
@@ 5 A* a = (A*)malloc(8); //o1
16void storePtr(void** ptrAddr) {
17 // look up the pointer pointing to object
18 // store the pointer address to the global data structure
19
@@
Figure 2. A Motivation Example.
Figure 3. Overview of LightDE.
To further illustrate the issue that existing methods of eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020) are not lightweight, we use the example in Figure 2 to explain this problem. In this example, a type $A$ object $o1$ is created on line 5, and the memory address of $o1$ is stored in the variable $A*\ a$. On line 7, an integer object $o2$ is allocated on the heap, and the memory address of $o2$ is stored in the variable $int*\ b$. On line 9, the memory address of $o2$ is assigned to the first field of object $o1$. The code segment highlighted in red in Figure 2 is an abstraction of the inserted instructions in existing dangling pointer elimination methods. The function $storePtr$ in Figure 2 also abstracts the insertion function from existing dangling pointer elimination methods. The primary responsibility of this function is to locate the object pointed to by a pointer and record the pointer’s address in the designated global data structure. It is worth noting that various methods may have slight differences in the specific operations of this function, but the basic functionality is similar. That is, tracking the pointers and the objects they point to in order to eliminate the corresponding dangling pointers when memory is freed. When the program executes the code in line 6 of Figure 2, the $storePtr$ function locates the object pointed to by $A*\ a$ and records the memory address of $A*\ a$ in the pointer set of the corresponding object. In this example, the pointer $A*\ a$ will be recorded in the pointer set of the object $o1$. Similarly, the code on lines 8 and 10 will trigger similar processes. This design means that whenever the program encounters a pointer assignment operation, it will perform object lookup and pointer memory address recording operations. This severely impacts the program’s performance. Through the example in Figure 2, it is easy to understand why existing dangling pointer elimination methods are not lightweight.
3. Overview of LightDE
We illustrate the working process of LightDE in Figure 3. First, we compile the program’s source code into LLVM Intermediate Representation (LLVM IR) using the LLVM compiler. We then perform pointer analysis based on the LLVM IR. Our structure-sensitive pointer analysis (Balatsouras and Smaragdakis, 2016; Barbar et al., 2020; Pearce et al., 2007; Avots et al., 2005) is conducted in two phases. The primary purpose of the first phase of pointer analysis is to identify the objects pointed to by pointers and the structural information of these objects. The second phase of the pointer analysis incorporates the structural information of objects obtained in the first phase and uses this structural information to filter out spurious pointers pointing to object fields. Next, based on the results of the pointer analysis, we categorize pointers into three types: heap pointers, global pointers, and stack pointers. This classification is essential because the elimination method for each type of pointer is different. Based on the results of the pointer analysis, we embed the pointer points-to information in the program’s data segment. However, it is not possible to determine the runtime memory address of a pointer in the static case. We will explain our method for storing pointer points-to relationships at compile time and mapping the results of static pointer analysis to runtime information in Section 4.2.
In the pointer analysis phase, our method of abstracting objects is to treat each heap allocation point as an object. This abstraction method is too coarse for eliminating dangling pointers, so we need to accurately identify each object allocated and released by the program. To address this issue, we use a multithreaded hash table (Gao et al., 2005) to map static objects to runtime objects. We insert a function after each object allocation point in the program to store the allocated object information and the static information of the object in our multithreaded hash table. Similarly, we insert a function before each object release point to identify the released objects and pass their information to the LightDE thread. We remove the program’s original memory release function, allowing the LightDE thread to handle memory space release to prevent UAF vulnerabilities during the elimination of dangling pointers by the LightDE thread. The creation of the LightDE thread occurs when the program starts execution. The function of this thread is to eliminate the corresponding dangling pointers based on the pointing relationships stored in the program’s data segment after receiving information about released objects passed from the program’s original thread. Finally, the LLVM IR with the inserted data and functions is compiled into a binary program.
4. Design
In this section, we first explain LightDE’s structure-sensitive pointer analysis method to obtain the pointing relationships of pointers and the structural information of objects. Then, we explain how LightDE maps the pointer analysis results to runtime information to accurately eliminate dangling pointers.
4.1. Structure-sensitive Pointer Analysis
In this section, we first explain our method for abstracting programs, then describe our pointer analysis method.
4.1.1. Program Abstraction

### v4

# LightDE: A Lightweight Method for Eliminating Dangling Pointers

@@
Abstract
The widespread presence of Use-After-Free (UAF) vulnerabilities poses a serious threat to software security, with dangling pointers being considered the primary cause of these vulnerabilities. However, existing methods for defending against UAF vulnerabilities by eliminating dangling pointers need to interrupt the program’s execution when encountering pointer assignment operations in order to store the memory addresses of the pointers in a specific data structure. This makes these methods not lightweight. To overcome this drawback, we propose a novel approach called LightDE. This method does not require storing the memory addresses of pointers during program execution. LightDE uses our proposed structure-sensitive pointer analysis method to determine which objects pointers point to and stores the pointing relationships in the program’s data segment during program compilation. Since LightDE only needs to verify if pointers identified by the pointer analysis point to released objects when eliminating dangling pointers, it is very lightweight. Our experimental results show that LightDE can effectively defend against UAF vulnerabilities and the performance overhead it introduces is very low.
Index Terms:
@@ Abstract
I Introduction
@@
Use-After-Free (UAF) vulnerabilities pose a serious threat to software security as they can lead to arbitrary code execution, privilege escalation, and data leakage [1, 2, 3]. Unfortunately, the number of UAF vulnerabilities has been increasing over the years, and their impact has become increasingly severe [1, 4, 5, 6]. Table 1 shows the changes in the number of UAF vulnerabilities in the NVD database from 2011 to 2023, as well as the number of high-threat UAF vulnerabilities. There are two types of UAF vulnerabilities: stack UAF and heap UAF [7, 8, 3, 9]. Stack UAF occurs when a dangling pointer is used to access stack memory that has already been freed. However, modern compilers generate warning messages for such vulnerabilities, making it easy for programmers to detect and fix them. Additionally, stack space allocation and deallocation are automatic, making stack UAF vulnerabilities difficult to exploit. Heap UAF refers to accessing freed heap memory through a dangling pointer. The difficulty in defending against heap UAF primarily stems from the flexible allocation and deallocation of heap memory, as well as the complexity of pointer propagation. All UAF vulnerabilities discussed in this paper are heap UAF.
Figure 1: Use-After-Free Vulnerabilities in NVD [10].
@@
Researchers have proposed various methods to defend against UAF vulnerabilities, but each approach has its limitations and drawbacks. One method involves checking the validity of pointer access during pointer dereferencing, which requires verifying memory access legality before each pointer dereference [11, 12, 13, 14, 15]. However, this method can lead to incompatibility with existing libraries and introduce unacceptable runtime overhead. An alternative approach modifies memory allocation and deallocation strategies, making it harder for attackers to exploit UAF vulnerabilities [16, 17, 18, 19, 20, 21]. However, this method only provides probabilistic defense, and attackers can bypass it through techniques like heap spraying and other attack methods. Recent research has proposed defending against UAF vulnerabilities by eliminating dangling pointers [22, 7, 23, 24, 25]. Eliminating dangling pointers is a very promising method for mitigating UAF vulnerabilities, as dangling pointers are the root cause of these vulnerabilities. However, existing methods for eliminating dangling pointers need to record the memory addresses of pointers during program execution. This makes these methods not lightweight.
@@
Since dangling pointers are the root cause of UAF vulnerabilities [22, 7, 23, 24, 25], this paper focuses on defending against UAF vulnerabilities by eliminating dangling pointers. Current methods [22, 7, 23, 24, 25] for defending against UAF vulnerabilities by eliminating dangling pointers rely on inserting functions after pointer assignment instructions during compilation to track pointers. This method requires the program to pause its normal execution flow when encountering a pointer assignment instruction during runtime, and to execute the logic of the inserted function. The purpose of inserting functions is to record the memory addresses of pointers. Some methods [22, 7, 23, 25] also need to locate the objects pointed to by the pointers. Although existing dangling pointer elimination schemes have designed various data structures to store the memory addresses of pointers to improve storage efficiency, their fundamental principles remain the same. In this process, recording the memory addresses of pointers and locating the objects they point to is time-consuming, which makes current dangling pointer elimination methods not lightweight.
@@
To address the issue of existing methods being unable to eliminate dangling pointers in a lightweight manner, we propose LightDE, a lightweight approach to eliminate dangling pointers to defend against UAF vulnerabilities. LightDE is automatically deployed during program compilation and is completely transparent to the programmer. LightDE’s uniqueness lies in its combination of our proposed structure-sensitive pointer analysis [26, 27, 28, 29] to determine which objects pointers reference at compile time, and storing these relationships in the program’s data segment during compilation. The significant advantage of our method over existing methods is that it does not require storing the memory addresses of pointers or locating the objects pointed to by the pointers during program execution. When an object is released, LightDE checks whether the pointer identified by the pointer analysis still points to the released object based on the pointer relationships stored in the program’s data segment, and eliminates the corresponding dangling pointers.
The main challenge in implementing LightDE lies in identifying the objects pointed to by pointers at compile time. Field-sensitive pointer analysis methods can analyze the objects pointed to by pointer fields. However, field-sensitive pointer analysis results do not include the type information of the objects. When it is unclear which specific field of an object points to another object, field-sensitive analysis cannot infer which pointer fields within the object might point to that object. As a result, it cannot completely eliminate dangling pointers within the object. Structure-sensitive pointer analysis results include the type information of objects. This method can use the object’s type information to infer the pointer fields within the object and their targets. But existing structure-sensitive pointer analysis cannot correctly analyze the case where an object has multiple types [34, 30] or cannot be used to analyze multi-threaded applications [31]. To address this problem, we propose a structure-sensitive, flow-insensitive pointer analysis method based on Andersen’s pointer analysis [26, 27, 28, 29] to acquire type information of objects. Our structure-sensitive pointer analysis is flow-insensitive, so our pointer analysis method can also be used to analyze pointer relationships in multithreaded applications. Although our pointer analysis method still has cases where it cannot identify the targets of pointer fields within objects, the results of our pointer analysis include the type information of objects. Therefore, when eliminating dangling pointers, we can check all the pointer fields of an object based on its type information to see if they point to the freed object. In Section IV-A1, we will illustrate with an example that proposing a completely new structure-sensitive pointer analysis method is necessary for eliminating dangling pointers.
@@
We implemented a prototype system of LightDE and demonstrated its effectiveness using real UAF vulnerabilities. Our experimental results on the SPEC CPU2006 [32] benchmark suite demonstrate that LightDE introduces very low runtime overhead, with a geometric mean runtime overhead of only 8.66%. The geometric mean of the memory overhead introduced by LightDE is 7.26%. Experimental results on the SPLASH-2X [33] benchmark suite show that LightDE scales well on multi-threaded applications. pSweeper also runs a separate thread on idle CPU cores to eliminate dangling pointers, but the geometric mean of the runtime overhead introduced by pSweeper is 17.2% which is significantly higher than the runtime overhead introduced by LightDE. Therefore, the reduced runtime overhead of LightDE is primarily due to determining and storing pointer relationships at compile time.
@@
1. We propose a lightweight method for eliminating dangling pointers, LightDE. Our approach determines pointer points-to relationships during program compilation and embeds this information into the program’s data segment at compile time. This allows our method to avoid interrupting the program’s execution flow during runtime to store the memory addresses of pointers or locate pointer targets, making LightDE very lightweight.
2. We propose a structure-sensitive pointer analysis method based on Andersen’s pointer analysis to obtain object structure information and improve the accuracy of pointer analysis results.
3. We developed a prototype system of LightDE based on LLVM-13 and SVF to validate our design. We conducted experiments to evaluate the LightDE prototype. The experimental results show that LightDE can effectively defend against UAF vulnerabilities while introducing very low performance overhead.
1 typedef struct {
@@ 5 A* a = (A*)malloc(8); //o1
16 void storePtr(void** ptrAddr) {
17 // store the pointer address
18
19 }
Figure 2: A Motivation Example.
Figure 3: The Compile-Time Workflow of LightDE.
II Motivation Example
To further illustrate the issue that existing methods of eliminating dangling pointers [22, 7, 23, 24, 25] are not lightweight, we use the example in Figure 2 to explain this problem. In this example, a type $A$ object $o1$ is created on line 5, and the memory address of $o1$ is stored in the pointer variable $A*\ a$. On line 7, an integer object $o2$ is allocated on the heap, and the memory address of $o2$ is stored in the pointer variable $int*\ b$. On line 9, the memory address of $o2$ is assigned to the first field of object $o1$. The code segment highlighted in red in Figure 2 is an abstraction of the inserted instructions from the existing dangling pointer elimination methods. The function $storePtr$ in Figure 2 is also abstracted from the existing dangling pointer elimination methods. The primary responsibility of this function is to record the pointer’s address in the designated global data structure. It is worth noting that various methods may have slight differences in the specific operations of this function, but the basic functionality is similar. That is, tracking the pointers in order to eliminate the corresponding dangling pointers when memory is freed. When the program executes the code on line 6 of Figure 2, the $storePtr$ function records the memory address of $A*\ a$. Similarly, the code on lines 8 and 10 will trigger similar processes. This design means that whenever the program encounters a pointer assignment operation, it records the pointer’s memory address. This process severely impacts the program’s performance. Through the example in Figure 2, it is easy to understand why existing dangling pointer elimination methods are not lightweight.
III Overview of LightDE
We use Figure 3 to illustrate the workflow and content insertion of LightDE during compilation, and use Figure 4 to illustrate the working state of LightDE during program execution. The workflow of LightDE at compile time can be divided into five parts: generating LLVM IR, performing pointer analysis, organizing the pointer analysis results, inserting the pointer analysis results and corresponding instructions into the LLVM IR, and recompiling the LLVM IR to generate the executable file. First, we compile the program’s source code into LLVM Intermediate Representation (LLVM IR) using the LLVM compiler. LightDE performs pointer analysis based on the LLVM IR. Our structure-sensitive pointer analysis [30, 34, 35, 36] is conducted in two phases. The primary purpose of the first phase of pointer analysis is to obtain the structural information of objects. The second phase of pointer analysis uses the object structure information obtained in the first phase to filter out spurious pointing relationships. Next, based on the results of pointer analysis, LightDE categorizes pointers into three types based on the objects they point to: heap pointers, global pointers, and stack pointers. Based on the results of the pointer analysis and the pointer types, LightDE embeds the pointer pointing information in the program’s data segment. However, it is not possible to determine the runtime memory address of a pointer in the static case. We will explain our method for storing pointer pointing relationships at compile time and mapping the results of static pointer analysis to runtime information in Section IV-B.
In the pointer analysis phase, our method of abstracting objects is to treat each heap allocation point as an object. This abstraction method is too coarse for eliminating dangling pointers, so we need to accurately identify every object that the program allocates and deallocates at runtime. To address this issue, LightDE uses a multi-threaded hash table [37] to map static objects to runtime objects. LightDE inserts a function after each object allocation point in the program to store the allocated object information and the static information of the object in our multi-threaded hash table. Similarly, LightDE inserts a function before each object release point to obtain the released objects from the hash table and pass their information to the LightDE thread. LightDE removes the program’s original memory release function, allowing the LightDE thread to handle memory space release to prevent UAF vulnerabilities during the elimination of dangling pointers. The creation of the LightDE thread occurs when the program starts executing. This thread has two functions. The first function is to organize all runtime objects created by the same allocation point together after receiving the object information from the program’s original thread. This organization facilitates the lookup of corresponding runtime objects based on static object information when eliminating dangling pointers. The second function is to eliminate corresponding dangling pointers based on the pointing relationships stored in the program’s data segment after receiving the information about released objects from the program’s original thread. Finally, the LLVM IR with the inserted data and functions is compiled into a binary program.
Figure 4: The Runtime Workflow of LightDE.
Figure 3 only illustrates the workflow of LightDE at compile time, but it does not reflect the operation of LightDE at runtime. Therefore, we use Figure 4 to explain the operation of LightDE at runtime. When the program starts executing, LightDE initializes our data structures and creates the LightDE thread. When the program executes to the object allocation instruction, the next instruction after the object allocation is our inserted function for tracking runtime objects. The function’s role is to store the starting memory address of the runtime object, the size of the object’s memory, and related static object information into a global multithreaded hash table. Additionally, this function stores this information in a multithreaded lock-free queue. Our thread retrieves object information from this lock-free queue and organizes all runtime objects created at the same allocation point together. When the program executes to the deallocation point for the object, the original deallocation function of the program has already been removed by LightDE at compile time. Therefore, our object deallocation management function is executed instead. The role of our object deallocation management function is to remove the deallocated object’s information from the global multithreaded hash table and store the deallocated object’s information into a lock-free queue shared with the LightDE thread. The LightDE thread obtains the object information to be released from this queue, and check whether the relevant pointer points to the object to be released based on the pointing relationship pre-stored in the program data segment. In Figure 4, we only show the entire process of allocating and releasing an object for one thread. However, the basic process for multiple objects and multiple threads is similar to the process shown in Figure 4.
IV Design
In this section, we first explain LightDE’s structure-sensitive pointer analysis method to obtain the pointing relationships of pointers and the structural information of objects. Then, we explain how LightDE maps static pointer analysis results to runtime information, enabling accurate elimination of dangling pointers.
IV-A Structure-sensitive Pointer Analysis
In this section, we first demonstrate through examples that our structure sensitivity is indispensable for eliminating dangling pointers. Then we explain our method for abstracting programs, and finally, we describe our pointer analysis method.
IV-A1 Motivating Example For Our Structure-Sensitivity
1typedef struct{
2 void* pf;
3} T1;
4typedef struct{
5 void *ps;
6 void* px;
7} T2;
8void main() {
9 int n;
10 T1* p = (T1*)malloc(16);//o1
11
12 T2* q = (T2*)p;
13 q->px = malloc(8); //o2
14 void* ptr = malloc(16); //o3
15 scanf(”%d”,&n); //n = 0 or 8
16 void* p1 =(void*) p + n;
17 *((void**)p1) = ptr;
18
19 free(ptr); //o3
@@ 12 T2* q = (T2*)p;
21}
Figure 5: A Motivating Example for the Necessity of Our Structure-Sensitivity.
We illustrate the innovation, necessity, and role of our structure-sensitive pointer analysis through the example in Figure 5. In the code shown in Figure 5, an object $o1$ is allocated on line 10 and cast to type $T1$, followed by a series of insignificant operations. On line 12, the code casts object $o1$ to type $T2$. On line 13, the code stores the memory address of object $o2$ into the $px$ field of object $o1$ through type $T2$. Existing structure-sensitive pointer analysis methods [34, 30] assume that an object has only one type throughout its entire lifecycle. Therefore, in this example, these existing methods assume that object $o1$ has only type $T1$. As a result, at line 13, they incorrectly conclude that object $o1$ does not have a $px$ field, leading to the inference that object $o1$ does not point to object $o2$. Additionally, the recent structure-sensitive pointer analysis method [31] cannot be used to analyze multi-threaded applications. The code on line 16 in the figure 5 modifies the pointer $p$ using the integer $n$, and the pointer $p1$ may point to the first field of object $o1$ or the second field of object $o1$. Therefore, it is impossible to ascertain which field of object $o1$ points to object $o3$ from analyzing the code at line 17. Field-sensitive pointer analysis will infer that object $o1$ points to object $o3$. However, since the field-sensitive pointer analysis results do not include the type information of the objects, it cannot infer the possible field offsets of object $o1$ that point to $o3$. This leads to an inability to eliminate the dangling pointer in object $o1$ after object $o3$ is freed on line 19. Similarly, our structure-sensitive pointer analysis results also cannot accurately determine which field of object $o1$ points to object $o3$. However, our structure-sensitive pointer analysis results include the type information $T1$ and $T2$ of object $o1$. Based on the type information of the object, our method can infer that either the field of object $o1$ at offset 0 points to object $o3$, or the field of object $o1$ at offset 8 points to object $o3$. After object $o3$ is released on line 19, LightDE can check the pointer fields at offsets 0 and 8 of object $o1$ to see if they point to object $o3$. The example in Figure 5 demonstrates that proposing a structure-sensitive pointer analysis method suitable for multithreading and multiple struct objects is essential for eliminating dangling pointers.
IV-A2 Program Abstraction